A large 35 mm thick slab of material has surface temperatures of 318.8 degrees K and 303.2 degrees K, and heat flux through the material of 35.1 W/m^2. Calculate the thermal conductivity of the slab.
Answer #1

Flux is given by:

Flux = k*dT/d where :

k = thermal conductivity

dT = temperature difference

d = thickness

Rearranging in terms of k:

k = flux*d/(dT)

= (35.1)(0.035m)/(318.8-303.2)

= 0.07875 W/(mK)